# Weakly Supervised Co-training with Swapping Assignments for Semantic Segmentation

### v1

@@ 3.2 Co-training with Swapping Assignments
Reliability based Adaptive Weighting. Segmentation performance depends heavily on the pseudo-labels. Despite the high-quality CPL generated by our guided CAMs, their reliability must be assessed, particularly in the initial training phases. Existing methods use post-refinement to increase reliability [70, 3]. As CoSA can generate online pseudo-labels, we propose to leverage confidence information to compensate the CAM2Seg loss during training.
Specifically, we propose to assess the perplexity scores for each pixel in $\hat{\mathcal{Y}}^{\text{CPL}}$ and leverage these scores to re-weight $\mathcal{L}_{\text{c2s}}$ for penalizing unreliable regions. However, estimating per-pixel perplexity is non-trivial. Through empirical analysis, we observe a noteworthy association between the confidence values of CAMs and their accuracy at each time-step. This correlation suggests that regions with extremely low or high confidence exhibit higher accuracy throughout training, as shown in Fig. 3(a)(b). To quantitatively model perplexity, we make two assumptions: i) the reliability of pseudo-labels is positively correlated with their accuracy, and ii) the perplexity score is negatively correlated with the reliability. Then, per-pixel perplexity of $\hat{\mathcal{Y}}^{\text{CPL}}_{x,y}$ is defined as:
@@
where the term within the logarithm denotes the normalized distance to $\xi$ in $[0,1]$. The logarithm ensures $\mathcal{P}_{x,y}\!\rightarrow\!+\infty$ as distance $\!\rightarrow\!0$, and $\mathcal{P}_{x,y}\!\rightarrow\!0$ as distance $\rightarrow\!1$. $\lambda_{\alpha}\in\mathbb{R}^{+}$ controls the perplexity score’s minimum value and $\lambda_{\beta}\in\mathbb{R}^{+}$ determines the sharpness or smoothness of the distribution. Higher $\mathcal{P}_{x,y}$ indicates confidence of $\hat{\mathcal{Y}}^{\text{CPL}}_{x,y}$ closer to threshold $\xi$. This observation is substantiated by Fig. 3 (a)(b), where confidence values near $\xi\!=\!0.5$ exhibit lower reliability. Furthermore, the correlation between perplexity and accuracy remains significant across various training time-steps and datasets, as depicted in Fig. 3(c)(d).
Since we hypothesize negative reliability-perplexity correlation, the reliability score can be defined as the reciprocal of perplexity. To accommodate reliability variation for different input, we use the normalized reliability as the per-pixel weights for $\mathcal{L}_{\text{c2s}}$. Thus, Reliability based Adaptive Weighting (RAW) is defined as: $\mathcal{W}_{x,y}^{\text{raw}}=|\mathcal{R}|/\sum_{i,j\in\mathcal{R}}{(\mathcal{P}_{i,j}\mathcal{P}_{x,y})}^{-1}$, where $|\mathcal{R}|$ represents total number of pixels in a batch. Then, the re-weighted CAM2seg loss for each position $(x,y)$ can be defined as
$\small\mathcal{L}_{\text{c2s}}(x,y)\!\!=\!\!-\mathcal{W}_{x,y}^{\text{raw}}\!\sum_{c=0}^{C}\!\Bigg{[}\!\mathbbm{1}\!\big{[}\hat{\mathcal{Y}}^{\text{CPL}}_{x,y}\!\!=\!\!c\big{]}\!\!\log\!\!\Bigg{(}\!\frac{\exp{\mathcal{S}^{c}_{x,y}}}{\sum_{k=0}^{C}\exp\mathcal{S}^{k}_{x,y}}\!\!\Bigg{)}\!\!\Bigg{]}.$ (5)
Dynamic Threshold. Existing WSSS work [52, 53] prescribes a fixed threshold to separate foreground and background, which neglects inherent variability due to prediction confidence fluctuation during training. Obviously, applying a fixed threshold in Fig. 3(a)(b) is sub-optimal. To alleviate this, we introduce dynamic thresholding. As shown in Fig. 3(e)(f), the confidence distribution reveals discernible clusters. We assume the foreground and background pixels satisfy a bimodal Gaussian Mixture distribution. Then, the optimal dynamic threshold $\xi^{\star}$ is determined by maximizing the Gaussian Mixture likelihood:
@@
$+$ $\prod_{x\in\{\mathcal{M}^{\prime}<\xi\}}\tilde{\pi}_{bg}\mathcal{N}\left(x\mid\tilde{\mu}_{bg},\tilde{\Sigma}_{bg}\right)~{},$
where $\mathcal{N}(x\mid\mu,\Sigma)$ denotes the Gaussian function and $\pi$, $\mu$, $\Sigma$ represent the weight, mean and covariance. To avoid mini-batch bias, we maintain a queue to fit GMM, with the current $\mathcal{M}^{\prime}$ batch enqueued and the oldest dequeued. This practice facilitates the establishment of a gradually evolving threshold, contributing to learning stabilization.
3.4 CAM Optimization.
@@
$\small\mathcal{S}^{\prime~{}c}_{x,y}=\left\{\begin{aligned} &-\infty,\!\!\!&\text{if $Y_{\text{gt}}^{c}=0$,}\\ &\mathcal{S}^{\prime~{}c}_{x,y},\!\!\!&\text{if $Y_{\text{gt}}^{c}\neq 0$,}\\ \end{aligned}\right.\;\;\;\;\hat{\mathcal{Y}}^{\text{SPL}}_{x,y}=\mathtt{Softmax}(\frac{\mathcal{S}^{\prime}_{x,y}}{\tau})~{},\vspace{-5pt}$ (7)
@@
where $\tau$ represents the softmax temperature to sharpen the $\hat{\mathcal{Y}}^{\text{SPL}}$. Then, we arrive Seg2CAM learning objective:
$\mathcal{L}_{\text{s2c}}=-\frac{1}{C|\mathcal{R}|}$ $\sum_{c=1}^{C}\sum_{x,y\in\mathcal{R}}\bigg{[}\hat{\mathcal{Y}}^{\text{SPL}}_{x,y}[c]\log(\sigma(\mathcal{M}^{c}_{x,y}))$ (8)
$+$ $(1-\hat{\mathcal{Y}}^{\text{SPL}}_{x,y}[c])\log(1-\sigma(\mathcal{M}^{c}_{x,y}))\bigg{]}~{},$
where $\mathcal{R}$ represents all the positions in SPL.
@@
Coexistence Problem in CAMs. Certain class activations often exhibit large false positive regions, where objects are incorrectly merged with surroundings, as shown in Fig. 8 (1st row). For instance, the classes ’bird’ and ’tree branches’, ’train’ and ’railways’, etc. frequently appear together in VOC. We refer to this issue as the coexistence problem. We hypothesize that the coexistence problem is attributed to three factors: i) Objects that coexist in images, such as ’tree branches’, are not annotated w.r.t. weak labels, which makes it challenging for a model to semantically distinguish coexistence. ii) Training datasets lack sufficient samples for such classes. iii) High-level feature maps, though rich in abstract representations and semantic information, lack essential low-level features such as edges, textures, and colors [24]. Thus, CAMs generated from the last layer are poor in low-level information for segmenting objects. Conversely, segmenting objects with high-level semantics is hindered due to factors i) and ii).
Figure 4: $\mathcal{M}$ and $\mathcal{M}^{\dagger}$ Comparisons. (a): mIoU vs. time-steps for $\mathcal{M}$ and $\mathcal{M}^{\dagger}$ on VOC val. (b): same as (a) but filtered by perplexity. (c): cases of coexistence issue in $\mathcal{M}$ but not in $\mathcal{M}^{\dagger}$.
@@
Contrastive Separation in CAMs. We posit that the effective usage of low-level information can alleviate the coexistence problem. Since shallower-layer feature maps contain more low-level information, we propose to extract CAMs from earlier in the backbone, denoted $\mathcal{M}^{\dagger}$, and present a comparison with $\mathcal{M}$ in Fig. 4, showing that substituting $\mathcal{M}$ with $\mathcal{M}^{\dagger}$ is not feasible due to the lower mIoU upperbound of $\mathcal{M}^{\dagger}$. If we consider the confident regions in $\mathcal{M}$ and $\mathcal{M}^{\dagger}$, i.e. filter by a low-pass perplexity $\mathcal{P}$, $\{\mathcal{M}^{\dagger}_{x,y}\!\mid\!\mathcal{P}_{x,y}\!\leq\!\epsilon\}$ are better than $\{\mathcal{M}_{x,y}\!\mid\!\mathcal{P}_{x,y}\!\leq\!\epsilon\}$, as shown in Fig. 4(b), where $\epsilon$ represents a low-pass coefficient. Fig. 4(c) illustrates the presence of coexistence issues in $\mathcal{M}$ but absence in $\mathcal{M}^{\dagger}$. Those findings suggest that $\mathcal{M}^{\dagger}$ is worse than $\mathcal{M}$ in general, but better for those regions with low perplexity.
In CoSA, we propose to regularize $\mathcal{M}$ by $\mathcal{M}^{\dagger\prime}$ (from AN). We define positive $\mathcal{R}^{+}_{i,j}$ and negative $\mathcal{R}^{-}_{i,j}$ regions as
@@
where $(i,j)\!\in\!\Omega$, $\Omega\!=\!\{(x,y)\!\mid\!\mathcal{P}_{x,y}\!\leq\!\epsilon\}$ is low-perplexity region in $\mathcal{M}^{\dagger\prime}$, and $\hat{y}^{\text{CPL}}$ represents the CPL of $\mathcal{M}^{\dagger\prime}$. Then, we arrive the contrastive separation loss for $\mathcal{M}$:
@@
where $L_{x,y}^{i,j}\!=\!\exp(l_{d}(\mathcal{M}_{i,j},\mathcal{M}_{x,y})/\tau)$, $l_{d}(a,b)$ measures the (a,b) distance, $\tau$ denotes the InfoNCE loss [47] temperature, and $K_{n,m}^{i,j}\!=\!\sum_{n,m\in\mathcal{R}^{-}_{i,j}}L_{n,m}^{i,j}$.
@@ 4.1 Experiment Details and Results
Datasets. We evaluate on two benchmarks: PASCAL VOC 2012 [19] and MS-COCO 2014 [41]. VOC encompasses 20 categories with train, val, and test splits of 1464, 1449, and 1456 images. Following WSSS practice [2, 3, 65], SBD [23] is used to augment the train split to 10,582. COCO contains 80 categories with train and val splits of approx. 82K and 40K images. Our model is trained and evaluated using only the image-level classification labels [^1] Not available for VOC test split and so not used in evaluation., and employing mIoU as evaluation metrics.
Implementation Details. Following [53], we use ImageNet pretrained ViT-base (ViT-B) [17] as the encoder. For classification, we use global max pooling (GMP) [51] and the CAM approach [72]. For the segmentation decoder, we use LargeFOV [10], as with [53]. ON is trained with AdamW [46]. The learning rate is set to 6E-5 in tandem with polynomial decay. AN is updated with a momentum of $0.9994$. For preprocessing, the images are cropped to $448^{2}$, then weak/strong augmentations are applied (see Supp. Materials). The perplexity constants $(\lambda_{\alpha},\lambda_{\beta})$ are set to $(0.8,1)$, GMM-fitting queue length is $100$, and softmax temperature $\tau$ is $0.01$. The low perplexity threshold $\epsilon$ is set to $1$ and the loss weight factors $(\lambda_{\text{c2s}},\lambda_{\text{s2c}},\lambda_{\text{csc}})$ to $(0.1,0.05,0.1)$.
Method Backbone train val
RRM [70] AAAI’2020 WR38 65.4
1Stage [3] CVPR’2020 WR38 66.9 65.3
AFA [52] CVPR’2022 MiT-B1 68.7 66.5
MCT [65] CVPR’2022 MCT 69.1
ViT-PCM [51] ECCV’2022 ViT-B 71.4 69.3
Xu et al. [66] CVPR’2023 ViT-B 66.3
ACR-ViT [31] CVPR’2023 ViT-B 70.9
CLIP-ES [42] CVPR’2023 ViT-B 75.0
ToCo [53] CVPR’2023 ViT-B 73.6 72.3
CoSA ViT-B 78.5 76.4
${}^{\bullet}$ ViT-B 78.9 77.2
Table 1: Comparisons of CPL. CAM pseudo-labels evaulation on VOC dataset. Backbone denotes the encoder used for generating the CAMs. $\bullet$ represents the ensemble of $\mathcal{M}^{\prime}$ and $\mathcal{M}^{\dagger\prime}$ in CoSA.
@@
CoSA-MS $\mathcal{I}$ SWIN 81.4 78.4${}^{\tiny{[3]}}$ 53.7
@@
CoSA $\mathcal{I}$ ViT-B 76.2 75.1${}^{\tiny{[4]}}$ 51.0
[…]
Table 2: Weakly Supervised Semantic Segmentation Results. Sup.: supervision type. Net.: segmentation backbone. $\mathcal{F}$: Fully supervised, $\mathcal{I}$: Image-level labels, $\mathcal{S}$: Saliency maps, $\mathcal{L}$: language models. *** represents CRF [10] postprocessing results.
CAM Quality Comparison. Tab. 1 shows CoSA CPL results on VOC compared with existing WSSS methods, using our $\hat{\mathcal{Y}}^{\text{CPL}}$ ($\xi\!\!=\!\!0.5$). Our method yields 78.5% and 76.4% mIoU on train and val. Notably, an ensemble of $\mathcal{M}^{\prime}$ and $\mathcal{M}^{\dagger\prime}$ improves performance to 78.9% and 77.2%, suggesting the activation of $\mathcal{M}^{\prime}$ is orthogonal to that of $\mathcal{M}^{\dagger\prime}$.
@@
Figure 5: Qualitative Comparison. The results are reported on the val splits of VOC (in R1 - R3) and COCO (in R4 - R6). The official codebases and provided weights for MCT [65], BECO [50], and ToCo [53] are used for this comparison. (best viewed under zoom; see Supp. Materials for more high-res Comparisons).

### v2

@@ 3.3 Segmentation Optimization
Figure 3: CPL Analysis (a): heatmap of CPL accuracy vs. confident ranges (x-axis) for different time-steps (y-axis) for VOC and COCO. (b): correlation between perplexity and accuracy of CPL for different time-steps. (c): distribution of CAMs’ confidence categorized by the proposed dynamic threshold on VOC. See Supp. for COCO analysis.
Reliability based Adaptive Weighting (RAW). Segmentation performance depends heavily on the reliability of the pseudo-labels. Thus, it is important to assess their reliability. Existing methods use post-refinement to enhance pseudo-label credibility [73, 3]. As CoSA can generate online CPL, we propose to leverage confidence information to compensate the CAM2Seg loss during training. Specifically, we propose to assess the perplexity scores for each pixel in $\hat{\mathcal{Y}}^{\text{CPL}}$ and leverage these scores to re-weight $\mathcal{L}_{\text{c2s}}$ for penalizing unreliable regions. However, estimating per-pixel perplexity is non-trivial. Through empirical analysis, we observe a noteworthy association between the confidence values of CAMs and their accuracy at each time-step. This correlation suggests that regions with extremely low or high confidence exhibit higher accuracy throughout training, as shown in Fig. 3(a). To quantitatively model perplexity, we make two assumptions: i) the reliability of pseudo-labels is positively correlated with their accuracy, and ii) the perplexity score is negatively correlated with the reliability. Then, per-pixel perplexity of $\hat{\mathcal{Y}}^{\text{CPL}}_{x,y}$ is defined as:
@@
where the term within the logarithm denotes the normalized distance to $\xi$ in $[0,1]$. The logarithm ensures $\mathcal{P}_{x,y}\!\rightarrow\!+\infty$ as distance $\!\rightarrow\!0$, and $\mathcal{P}_{x,y}\!\rightarrow\!0$ as distance $\rightarrow\!1$. $\lambda_{\alpha}\in\mathbb{R}^{+}$ controls the perplexity score’s minimum value and $\lambda_{\beta}\in\mathbb{R}^{+}$ determines the sharpness or smoothness of the distribution. Higher $\mathcal{P}_{x,y}$ indicates confidence of $\hat{\mathcal{Y}}^{\text{CPL}}_{x,y}$ closer to threshold $\xi$. This observation is substantiated by Fig. 3(a), where confidence values near $\xi\!=\!0.5$ exhibit lower reliability. Furthermore, the correlation between perplexity and accuracy remains significant across various training time-steps and datasets, as depicted in Fig. 3(b).
Since we hypothesize negative reliability-perplexity correlation, the reliability score can be defined as the reciprocal of perplexity. To accommodate reliability variation for different input, we use the normalized reliability as the per-pixel weights for $\mathcal{L}_{\text{c2s}}$. This arrives our RAW-based CAM2Seg objective:
$\small\mathcal{L}_{\text{c2s}}(x,y)\!\!=\!\!-\frac{|\mathcal{R}|}{\sum_{i,j\in\mathcal{R}}{(\mathcal{P}_{i,j}\mathcal{P}_{x,y})}^{-1}}\!\sum_{c=0}^{C}\!\Bigg{[}\!\mathbbm{1}\!\big{[}\hat{\mathcal{Y}}^{\text{CPL}}_{x,y}\!\!=\!\!c\big{]}\!\!\log\!\!\Bigg{(}\!\frac{\exp{\mathcal{S}^{c}_{x,y}}}{\sum_{k=0}^{C}\exp\mathcal{S}^{k}_{x,y}}\!\!\Bigg{)}\!\!\Bigg{]},$ (5)
where $|\mathcal{R}|$ represents total number of pixels in a batch.
Dynamic Threshold. Existing WSSS work [52, 53] prescribes a fixed threshold to separate foreground and background, which neglects inherent variability due to prediction confidence fluctuation during training. Obviously, applying a fixed threshold in Fig. 3(a) is sub-optimal.
To alleviate this, we introduce dynamic thresholding. As shown in Fig. 3(c), the confidence distribution reveals discernible clusters. We assume the foreground and background pixels satisfy a bimodal Gaussian Mixture distribution. Then, the optimal dynamic threshold $\xi^{\star}$ is determined by maximizing the Gaussian Mixture likelihood:
@@
where $\mathcal{N}(x|\mu,\Sigma)$ denotes the Gaussian function and $\pi$, $\mu$, $\Sigma$ are the weight, mean and covariance. To avoid mini-batch bias, we maintain a queue to fit GMM, with the current $\mathcal{M}^{\prime}$ batch enqueued and the oldest dequeued. This facilitates establishment of a gradually evolving threshold, contributing to learning stabilization.
3.4 CAM Optimization
Figure 4: Illustration of Coexistence Problem in CAMs. The first row shows the input images. The second row shows the coexistence problem e.g. ’bird’ with ’branches’, ’train’ with ’railways’ and ’boat’ with ’the sea’.
@@ 3.4 CAM Optimization
$\small\mathcal{S}^{\prime~{}c}_{x,y}=\left\{\begin{aligned} &-\infty,\!\!\!&\text{if $Y_{\text{gt}}^{c}=0$,}\\ &\mathcal{S}^{\prime~{}c}_{x,y},\!\!\!&\text{if $Y_{\text{gt}}^{c}\neq 0$,}\\ \end{aligned}\right.\;\;\;\;\hat{\mathcal{Y}}^{\text{SPL}}_{x,y}=\mathtt{Softmax}(\frac{\mathcal{S}^{\prime}_{x,y}}{\tau})~{},$ (7)
@@
where $\tau$ denotes the temperature to sharpen $\hat{\mathcal{Y}}^{\text{SPL}}_{x,y}$. Let $\mathcal{R}$ be all the positions in SPL, then the Seg2CAM learning objective is defined as:
$\small\mathcal{L}_{\text{s2c}}=-\frac{1}{C|\mathcal{R}|}\sum_{c=1}^{C}\sum_{x,y\in\mathcal{R}}\bigg{[}\hat{\mathcal{Y}}^{\text{SPL}}_{x,y}[c]\log(\sigma(\mathcal{M}^{c}_{x,y}))+(1-\hat{\mathcal{Y}}^{\text{SPL}}_{x,y}[c])\log(1-\sigma(\mathcal{M}^{c}_{x,y}))\bigg{]}~{}.$ (8)
@@
Coexistence Problem in CAMs. Certain class activations often exhibit large false positive regions, where objects are incorrectly merged with surroundings, as shown in Fig. 4. For instance, the classes ‘bird’ and ‘tree branches’, ‘train’ and ‘railways’, etc. frequently appear together in VOC dataset. We refer to this issue as the coexistence problem. We hypothesize that the coexistence problem is attributed to three factors: i) Objects that coexist in images, such as ‘tree branches’, are not annotated w.r.t. weak labels, which makes it challenging for a model to semantically distinguish coexistence. ii) Training datasets lack sufficient samples for such classes. iii) High-level feature maps, though rich in abstract representations and semantic information, lack essential low-level features such as edges, textures, and colors [24]. Thus, CAMs generated from the last layer are poor in low-level information for segmenting objects. Conversely, segmenting objects with high-level semantics is hindered due to factors i) and ii).
@@
Contrastive Separation in CAMs. We posit that the effective usage of low-level information can alleviate the coexistence problem. Since shallower-layer feature is rich in low-level info [72], we propose to extract CAMs $\mathcal{M}^{\dagger}$ from an earlier layer, and present its comparison with $\mathcal{M}$ in Fig. 5, showing that directly substituting $\mathcal{M}$ with $\mathcal{M}^{\dagger}$ is not feasible due to the lower mIoU upperbound of $\mathcal{M}^{\dagger}$. However, if we consider the confident regions in $\mathcal{M}$ and $\mathcal{M}^{\dagger}$, i.e. filter by a low-pass perplexity, then $\{\mathcal{M}^{\dagger}_{x,y}\!\mid\!\mathcal{P}_{x,y}\!\leq\!\epsilon\}$ result in higher mIoU than $\{\mathcal{M}_{x,y}\!\mid\!\mathcal{P}_{x,y}\!\leq\!\epsilon\}$, as shown in Fig. 5(b), where $\epsilon$ denotes a low-pass coefficient. Further, we observe in some examples the presence of coexistence issues in $\mathcal{M}$ but absence in $\mathcal{M}^{\dagger}$ as shown in Fig. 5(c). This suggests that $\mathcal{M}^{\dagger}$ performs worse than $\mathcal{M}$ in general, but better for those regions with low perplexity. Driven by these findings, we propose to regularize $\mathcal{M}$ by $\mathcal{M}^{\dagger\prime}$ (from AN). Specifically, $\mathcal{M}^{\dagger\prime}$ after a low-pass filter are used to determine the positive $\mathcal{R}^{+}_{i,j}$ and negative $\mathcal{R}^{-}_{i,j}$ regions:
@@
where $(i,j)\!\in\!\Omega$, $\Omega\!=\!\{(x,y)\!\mid\!\mathcal{P}_{x,y}\!\leq\!\epsilon\}$ is low-perplexity region in $\mathcal{M}^{\dagger\prime}$, and $\hat{y}^{\text{CPL}}$ represents the CPL of $\mathcal{M}^{\dagger\prime}$. Then, we have contrastive separation loss for $\mathcal{M}$:
@@
where $L_{x,y}^{i,j}\!=\!\exp(l_{d}(\mathcal{M}_{i,j},\mathcal{M}_{x,y})/\tau)$, $l_{d}(a,b)$ measures the similarity between (a,b), $\tau$ denotes the InfoNCE loss [47] temperature, and $K_{n,m}^{i,j}\!=\!\sum_{n,m\in\mathcal{R}^{-}_{i,j}}L_{n,m}^{i,j}$.
Figure 5: $\mathcal{M}$ and $\mathcal{M}^{\dagger}$ Comparisons. (a): mIoU vs. time-steps for $\mathcal{M}$ and $\mathcal{M}^{\dagger}$ on VOC val. (b): same as (a) but filtered by perplexity. (c): cases of coexistence in $\mathcal{M}$ but not in $\mathcal{M}^{\dagger}$.
@@ 4.1 Experiment Details and Results
Datasets. We evaluate on two benchmarks: VOC [19] and COCO [41]. VOC encompasses 20 categories with train, val, and test splits of 1464, 1449, and 1456 images. Following WSSS practice [2, 3, 65], SBD [23] is used to augment the train split to 10,582. COCO contains 80 categories with train and val splits of approx. 82K and 40K images. Our model is trained and evaluated using only the image-level classification labels [^1] Not available for VOC test split and so not used in evaluation., and employing mIoU as evaluation metrics.
Implementation Details. Following [53], we use ImageNet pretrained ViT-base (ViT-B) [17] as the encoder. For classification, we use global max pooling (GMP) [51] and the CAM approach [75]. For the segmentation decoder, we use LargeFOV [8], as with [53]. ON is trained with AdamW [46]. The learning rate is set to 6E-5 in tandem with polynomial decay. AN is updated with a momentum of $0.9994$. For preprocessing, the images are cropped to $448^{2}$, then weak/strong augmentations are applied (see Supp.). The perplexity constants $(\lambda_{\alpha},\lambda_{\beta})$ are set to $(0.8,1)$, GMM-fitting queue length is $100$, and softmax temperature $\tau$ is $0.01$. The low perplexity threshold $\epsilon$ is set to $1$ and the loss weight factors $(\lambda_{\text{c2s}},\lambda_{\text{s2c}},\lambda_{\text{csc}})$ to $(0.1,0.05,0.1)$.
@@
CoSA-MS $\mathcal{I}$ SWIN 81.4 78.4[3] 53.7
@@
SeCo [69] CVPR’2024 $\mathcal{I}$ ViT-B 74.0 73.8 46.7
CoSA $\mathcal{I}$ ViT-B 76.2 75.1[4] 51.0
CoSA $\mathcal{I}$ ViT-B 76.4 75.2[5] 51.1
Table 1: Weakly Supervised Semantic Segmentation Results. Sup.: supervision type. Net.: segmentation backbone. $\mathcal{F}$: Fully supervised, $\mathcal{I}$: Image-level labels, $\mathcal{S}$: Saliency maps, $\mathcal{L}$: language models. * represents CRF [8] postprocessing results.
CAM Quality Comparison. Tab. 2 shows CoSA’s CPL results compared with existing WSSS methods. Our method yields 78.5% and 76.4% mIoU on train and val. Notably, an ensemble of $\mathcal{M}^{\prime}$ and $\mathcal{M}^{\dagger\prime}$ improves performance to 78.9% and 77.2%, suggesting the activation of $\mathcal{M}^{\prime}$ is orthogonal to that of $\mathcal{M}^{\dagger\prime}$.
Method ViT-PCM [51] ACR-ViT [31] CLIP-ES [42] SeCo [69] ToCo [53] CoSA CoSA
train 71.4 70.9 75.0 76.5 73.6 78.5 78.9
val 69.3 72.3 76.4 77.2
Table 2: Comparisons of CPL. All methods use ViT as the backbone for generating the CAMs on VOC dataset. $\bullet$ represents the ensemble of $\mathcal{M}^{\prime}$ and $\mathcal{M}^{\dagger\prime}$ in CoSA.
@@
Figure 6: Qualitative Comparison. The results are reported on the val splits of VOC (in R1 - R3) and COCO (in R4 - R6). The official codebases and provided weights for MCT [65], BECO [50], and ToCo [53] are used for this comparison. (best viewed under zoom; see Supp. for more).
Figure 7: Effect of CSC. The class activation for bird, train, plane, boat and car are presented from top to bottom.
mIoU (inc.) Base. GC SA RAW CSC DT VOC COCO 55.96 37.32 63.09 (+7.13) 42.55 (+5.23) 64.41 (+8.45) 43.92 (+6.60) 68.22 (+12.26) 45.39 (+8.07) 71.66 (+15.70) 47.10 (+9.78) 75.54 (+19.58) 49.67 (+12.35) 76.19 (+20.23) 51.00 (+13.68)
mIoU (dec.) CoSA GC SA RAW CSC DT VOC COCO 76.19 51.00 75.54 (-0.65) 49.67 (-1.33) 69.89 (-6.30) 45.95 (-5.05) 72.45 (-3.74) 47.83 (-3.17) 72.10 (-4.09) 49.04 (-1.96) 74.12 (-2.07) 49.67 (-1.33)